# DRAGON: Dynamic RAG Benchmark On News

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) is a widely adopted approach for improving the factuality of large language models (LLMs) by incorporating external knowledge at inference time. Although there exist multiple RAG benchmarks for English, evaluation resources for other languages, including Russian, remain scarce and static, failing to capture the dynamic nature of real-world deployments.
In this work, we present DRAGON (Dynamic RAG Benchmark On News), the first dynamic benchmark for evaluating RAG systems in Russian on a changing news corpora. DRAGON is built upon a regularly updated corpus of Russian news and public documents and supports comprehensive evaluation of both the retriever and generator components. Question generation is performed automatically with the use of Knowledge Graph constructed from the corpus and enables the extraction of four core question types aligned with distinct subgraph patterns. We release a complete evaluation framework comprising the pipeline for automatic question generation, evaluation scripts, which are potentially reusable for other languages and multilingual settings, and benchmark data. We also launch a public leaderboard to encourage community participation and comparison.
DRAGON: Dynamic RAG Benchmark On News
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful instrument for enhancing the domain adaptation and factuality of large language models (LLMs) by incorporating external knowledge retrieved at inference time. This approach enables more up-to-date and grounded responses without the need for costly re-training. As RAG-based systems expand to applications such as open-domain QA, customer support, and enterprise search, their standardized evaluation becomes crucial for assessing reliability and utility.
While several RAG benchmarks have been proposed for the English language, there are significantly fewer evaluation resources for other languages, including Russian.
In this work, we focus on the evaluation of RAG systems for the Russian language. We introduce DRAGON: Dynamic RAG Benchmark On News (Fig. 1) a novel benchmark that reflects realistic usage patterns by leveraging a regularly updated knowledge base derived from current news sources. To foster transparency and community engagement, we publicly release an evaluation framework which comprises the codebase for automatic question generation, evaluation scripts, and a dynamic leaderboard to track progress on RAG-based systems in Russian. While the benchmark targets Russian, the framework is potentially extendable to other languages and multilingual scenarios, making it broadly applicable. Our contributions are as follows:
(i) We propose DRAGON111The video demonstration of the evaluation tool is available on YouTube., the first RAG benchmark with a regularly updated knowledge base based on news corpora for Russian, designed to evaluate RAG systems in a dynamic setup.
@@
(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/spaces/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
@@ 2 Related Work
Proper assessment of RAG systems requires holistic benchmarks that evaluate both the retrieval and the generation components. For English, a growing body of research has focused on designing such benchmarks across a range of reasoning and grounding challenges.
KILT Petroni et al. (2021) provides a unified benchmark over a fixed Wikipedia snapshot, with an emphasis on source attribution and retrieval grounding. Conversational settings with multi-turn dialogue scenarios are investigated in mtRAG Katsis et al. (2025) and RAD-Bench Kuo et al. (2025).
The factual consistency of generated responses in RAG has become a critical challenge. CRAG Yang et al. (2024) introduces a comprehensive benchmark based on five main RAG characteristics. RAGAS Es et al. (2024) proposes a reference-free evaluation suite that captures context relevance, faithfulness, and answer completeness, along with an open-source API for practical benchmarking.
Despite these advances, most of the RAG benchmarks focus on English. Evaluation for other languages remains limited. NoMIRACL Thakur et al. (2024) provides multilingual QA tasks in 18 languages and analyzes retrieval failures and hallucination rates. For Russian, RusBEIR Kovalev et al. (2025) offers retrieval-only evaluation, but a complete end-to-end RAG benchmark has yet to be developed.
@@
To address this need we present DRAGON – a dynamic, regularly updated benchmark for Russian-language RAG systems based on real-world, shifting corpora.
3 System Demo
@@
The benchmark is designed to evaluate RAG systems in a dynamically evolving news domain. Its architecture prioritizes modularity, automation, and reproducibility while addressing the core challenges Yu et al. (2024) in the RAG evaluation landscape such as temporal aspects of information, dynamic and vast knowledge sources, and factuality and faithfulness in generation.
Dynamic Benchmark Architecture
The entire pipeline of the benchmark architecture is shown in Fig. 2. Below, each step is described in more detail.
@@
Data Acquisition and Processing: We maintain a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx. A for the news sources list) on a daily basis. Newly parsed content is synchronized into our storage. To avoid redundancy and ensure incremental updates, a scheduled automated job identifies differences with the previous dataset revision and extracts updated segments for downstream processing.
@@
This design ensures the benchmark reflects evolving real-world distributions and mitigates the risks of overfitting to static datasets. The pipeline further ensures that newly surfaced topics and entities from the news stream are constantly incorporated into the benchmark.
@@
QA Dataset Formation: The process of creating the questions and answers based on the updated increment of news data is described in detail in Sec. 4. The pipeline transforms the generated QA pairs into several HF datasets, which form the core of our benchmark:
@@
All datasets are versioned and uploaded to Hugging Face with incrementally updated revision numbers. This versioning mechanism ensures reproducibility and provides users with stable snapshots for further experimentation.
@@
To facilitate seamless evaluation for users, we provide a PyPi-hosted Python library rag_bench, which offers an interface to:
Fetch the latest version of the public datasets by dynamically resolving the latest Hugging Face revision; Observe the RAG system baseline which can be adopted for the target one; Evaluate RAG system and package results for submission; Submit results via API to our evaluation portal; Calculate retrieval and generative metrics locally using the sandbox datasets.
@@
"model_answer": "Answer": "Top"
These results encode both the retrieved public_ids and the generated answers, decoupling the user’s model output from any private evaluation artifacts. This separation allows secure evaluation without exposing ground-truth data.
@@
A Hugging Face Gradio Space serves as the public Leaderboard (see Appx. B for a screenshot). Results are committed in a version-controlled results.json file, automatically updated by the validation portal upon approval.
To reduce latency and improve benchmarking coverage, we support automatic evaluation for selected pre-approved baselines, which include several popular LLMs and retrieval embedding models (see Appx. C for the details). The results are computed via the same rag_bench client.
@@ 4 Methodology
The Data Generation pipeline (Fig. 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation. KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph. The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with LLM.
Knowledge Graph Extraction
To enable greater control over question generation, we developed a Knowledge Graph Extraction module, inspired by Chepurova et al. (2024).
@@
At the first stage, the LLaMa 3.3 70B Instruct444https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Grattafiori et al. (2024) is used for the extraction of triplet-candidates. Each triplet consists of a head entity or subject, a relation, and a tail entity or object.
All the extracted entities are used to match with entities from Wikidata555We use Russian subgraph of Wikidata as a base one. Vrandečić and Krötzsch (2014). To do this, we search by API in this open knowledge base for the entities extracted in the previous step. Afterward, the found entities are vectorized using an embedding model and saved as a vector database. Then, for each entity extracted in the first step, we list five of the closest entity candidates in the vector representation base.
At the final stage, we use the same LLaMa 3.3 70B Instruct for normalization and unification of the extracted entities and relations. For each entity and relation in the triplet, the language model analyzes the list of candidates obtained at the previous stage and decides on the possibility of replacing the original entity or relation with one of the candidates. The triplet context is taken into account by adding an original text. This approach allows the unification of naming variants for the same entity, which is critical for the subsequent construction of the knowledge graph.
We aim to create a graph that contains novel information. To ensure this, we discard any triplets that are completely matched with saved entities and relations from the open knowledge base (Wikidata). We assume that facts not included in Wikidata are novel and thus unlikely to have been seen by the language model.
Question Generation
The question generation stage begins with subgraph extraction. We systematically sample all subgraphs that match one of the four predefined structural templates, each corresponding to one of the question types. We employ a typology of four question types, inspired by Yang et al. (2024): simple, set, multi-hop, and conditional questions (see Appx. D for a detailed description).
@@
A subgraph is passed to the language model, which is instructed to generate both a natural language question and a corresponding answer. The question is formulated as a coherent sentence, while the answer consists of one or more entities derived from the subgraph.
QA Filtering
The dataset requires filtering to improve the quality of question-answer pairs and mitigate language model errors. Firstly, the linguistic acceptability of each question is measured using a model trained on the RuCoLa dataset666https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola Mikhailov et al. (2022) model, which scores sentences based on their grammatical correctness and fluency.
Secondly, we extract Named Entities from the original document, solving the NER task with Natasha library. We look for these entities in generated question-answer pairs. The pairs that do not contain named entities are discarded to avoid examples relying on general knowledge. The questions are passed to the small LLMs (namely, we use Qwen 2.5 7B Instruct Team (2024), Llama 3 8B Grattafiori et al. (2024)) without context to detect the simplistic pairs that do not challenge the language model.
Thirdly, each generated pair is checked for compliance with the source subgraph. Either question or answer should contain an entity from the subgraph. To ensure this, we search the closest by Levenshtein distance Levenshtein et al. (1966) entity from the subgraph. We consider an entity to be found if the smallest distance from the entities in the subgraph is under the predefined threshold. We assume that the generated set should contain all the entities from the subgraph and not contain any extra ones. Otherwise, there is a generation error or the question type mismatch (see Appx. E).
The final stage uses the LLM-as-Judge approach to measure the question and answer quality. The pretrained judge POLLUX 7B Martynov et al. (2025) specifically trained for a fine-grained evaluation in Russian is adopted for this task. The model evaluates each sample against 8 generative criteria developed to assess the quality of the question-answer pair and its context dependence (the set description is given in Appx. F). As a result, questions of low quality are excluded from the final test set. The evaluation process is described in Appx. G. After the filtering stage 150 questions for each category form the final test dataset.
@@ 5 Experimental Setup
To construct our experimental RAG systems, we used the LangChain framework777https://pypi.org/project/langchain/. All texts from the Public Texts dataset are split into chunks of length 500 with an overlap of 100 characters. Each chunk is vectorized using the retrieval model of the evaluating RAG system with the corresponding document prefixes (see Appx. H.1 for their exact formulation), and the resulting vectors are stored in a vector database.
During the search phase, we use the prompted retrieval model to find five of the most relevant texts that match the user’s query. Retrieved chunks are incorporated into a prompt (see Appx. H.2) provided to the LLM of the evaluated RAG system. If the total length of the filled-in prompt exceeds the model’s maximum context length, the contextual information is truncated to the required size. To accelerate LLM inference, we utilize the vLLM framework888https://github.com/vllm-project/vllm Kwon et al. (2023). Other configuration details are described in Appx. H.3.
Scoring metrics
The performance of retrieval is measured with the 3 main metrics: Hit Rate, Recall, and NDCG. End-to-end RAG-system evaluation is performed via ROUGE-L, Substring Matching (SM), and Judge Score. See Appx. H.4 for their description.

### v3

# DRAGOn: Designing RAG On Periodically Updated Corpus

@@ Abstract
This paper introduces DRAGOn, method to design a RAG benchmark on a regularly updated corpus. It features recent reference datasets, a question generation framework, an automatic evaluation pipeline, and a public leaderboard. Specified reference datasets allow for uniform comparison of RAG systems, while newly generated dataset versions mitigate data leakage and ensure that all models are evaluated on unseen, comparable data. The pipeline for automatic question generation extracts the Knowledge Graph from the text corpus and produces multiple question-answer pairs utilizing modern LLM capabilities. A set of diverse LLM-as-Judge metrics is provided for a comprehensive model evaluation. We used Russian news outlets to form the datasets and demonstrate our methodology. We launch a public leaderboard to track the development of RAG systems and encourage community participation.
DRAGOn: Designing RAG On Periodically Updated Corpus
Fedor Chernogorskii2,1, Sergei Averkiev1, Liliya Kudraleeva3, Zaven Martirosian1,4,8, Maria Tikhonova1,5, Valentin Malykh6,3,7, Alena Fenogenova1,5 1SberAI, 2MBZUAI, 3ITMO, 4MISIS, 5HSE University, 6MWS AI, 7IITU, 8YSDA Correspondence: fechernogor@gmail.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful tool for enhancing the domain adaptation and factuality of large language models (LLMs) by incorporating external knowledge retrieved at inference time. This approach enables more up-to-date and grounded responses without the need for costly re-training. As RAG-based systems expand to applications such as open-domain QA, customer support, and enterprise search, their standardized evaluation remains a challenge. It may be unclear whether strong performance of a system is due to the quality of its retriever-generator pipeline or because the underlying LLM has been exposed to portions of the test data during training. It is possible that a static benchmark will become contaminated over time.
Several existing RAG evaluation frameworks Es et al. (2024); Lyu et al. (2025) provide pipelines for automatic generation of question-answer pairs, typically assuming that users will build their own evaluation datasets. While this enables domain-specific benchmarking, it is labor-intensive and difficult to maintain. Furthermore, when the retrieval corpus is continuously updated during deployment, results may become non-reproducible, as the model may face a different knowledge distribution than the one used during its initial evaluation.
In this work, we introduce DRAGOn: Designing RAG On Periodically Updated Corpus a novel methodology that reflects realistic usage patterns by leveraging a regularly updated knowledge base. We also release a benchmark built on current news sources using this methodology, which can be readily adapted to other document domains, such as scientific papers or court decisions. To foster transparency and community engagement, we publicly release an evaluation framework which comprises the codebase for automatic question generation, evaluation scripts, and a dynamic leaderboard to track progress on RAG-based systems in Russian. Although the benchmark targets Russian, the framework is potentially extendable to other languages and multilingual scenarios, making it broadly applicable. Our contributions are as follows:
(i) We propose DRAGOn111The video demonstration of the evaluation tool is available on YouTube. the methodology to develop a RAG benchmark with a regularly updated knowledge base, designed to evaluate RAG systems in a dynamic setup; we develop a benchmark on Russian news corpora as a reference for the proposed methodology.
@@
(iii) We launch a regularly updated public leaderboard333https://huggingface.co/spaces/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
@@ 2 Related Work
Evaluating retrieval-augmented generation (RAG) systems poses unique challenges, as it requires datasets that jointly assess both the retrieval and generation components. Constructing such benchmarks is costly and time-consuming because it involves curating large collections of text–question–answer triplets. To alleviate this, several works have explored synthetic data generation to automate question and answer creation Es et al. (2024); Lyu et al. (2025), often leveraging domain-specific pipelines or knowledge graphs for better control over content and difficulty.
Early RAG benchmarks such as KILT Petroni et al. (2021) unified multiple English-language datasets over a fixed Wikipedia snapshot, emphasizing source attribution and retrieval grounding. More recent efforts have extended the evaluation to multi-turn conversational and reasoning-intensive scenarios, as seen in mtRAG Katsis et al. (2025) and RAD-Bench Kuo et al. (2025). The CRAG benchmark Yang et al. (2024) further focuses on factual consistency, capturing five key aspects of RAG system behavior. Complementarily, RAGAS Es et al. (2024) provides a reference-free evaluation framework measuring context relevance, faithfulness, and answer completeness, and offers an open-source API for reproducible benchmarking.
Dynamic and time-sensitive evaluation has emerged as another important dimension. RealTime QA Kasai and et al. (2023) introduced a benchmark for evaluating systems on continuously evolving information sources, reflecting real-world deployment settings. The news domain, with its frequent updates and temporal drift, has thus become a popular testbed for such studies Tang and Yang (2024); Chen et al. (2024).
Despite these advances, the field still lacks a universal, contamination-free, and continuously updated benchmarks White et al. (2024). This gap hinders fair and reproducible comparison across RAG systems and motivates the development of dynamic, standardized evaluation resources.
@@
To address this need, we present DRAGOn – a methodology to develop dynamic, regularly updated benchmarks for RAG systems based on real-world, shifting corpora.
3 Benchmark Design
@@
Using our methodology, one can develop a benchmark designed to evaluate RAG systems in a dynamically evolving news domain. The benchmark’s architecture prioritizes modularity, automation, and reproducibility while addressing the core challenges Yu et al. (2024) in the RAG evaluation landscape, such as the temporal aspects of information, the vast and dynamic sources of knowledge, and the factuality and faithfulness in generation. The entire pipeline of the benchmark architecture is shown in Fig. 2. Below, each step is described in more detail.
@@
Data Acquisition and Processing: We maintain a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia on a daily basis. The parsed content is synchronized with our storage. To avoid redundancy and ensure incremental updates, a scheduled automated job identifies differences with the previous dataset revision and extracts updated segments for downstream processing.
@@
This design ensures that the benchmark reflects evolving real-world distributions and mitigates the risks of overfitting to static datasets. The pipeline further ensures that newly surfaced topics and entities from the news stream are constantly incorporated into the benchmark.
@@
QA Dataset Formation: The process of creating questions and answers based on the updated increment of the news data is described in detail in Sec. 4. The pipeline transforms the generated QA pairs into several HF datasets, which form the core of the benchmark:
@@
To facilitate seamless evaluation for users, we provide a PyPi-hosted Python library rag_bench, which offers an interface to: Fetch the latest version of the public datasets by dynamically resolving the latest Hugging Face revision; Observe the RAG system baseline, which can be adopted for the target one; Evaluate RAG system and package results for submission; Submit results via API to our evaluation portal; Calculate retrieval and generative metrics locally using the sandbox datasets.
@@
"model_answer": "Answer"
These results encode both the retrieved public_ids and the generated answers, decoupling the user’s model output from any private evaluation artifacts. This separation allows for secure evaluation without exposing ground-truth data.
@@
A Hugging Face Gradio Space serves as the public Leaderboard. The results are committed in a version-controlled results.json file, automatically updated by the validation portal upon approval.
To reduce latency and improve benchmarking coverage, we support automatic evaluation for selected pre-approved baselines, which include several popular LLMs and retrieval embedding models. The results are computed via the same rag_bench client.
@@ 4 Dataset Generation
The Data Generation pipeline (see Fig. 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation. KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph. The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with an LLM.
4.1 Knowledge Graph Extraction
To achieve fine-grained control over automatic question generation, we designed a Knowledge Graph (KG) Extraction module inspired by Chepurova et al. (2024). This component transforms unstructured news texts into a structured set of factual triplets that later guide question creation.
@@ 4.1 Knowledge Graph Extraction
We use LLaMa 3.3 70B Instruct444https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Grattafiori et al. (2024) to extract candidate factual triplets from the corpus. Each triplet has the form (head entity – relation – tail entity), corresponding to the subject, predicate, and object in the original sentence.
The extracted entities are matched with the Russian subgraph of Wikidata Vrandečić and Krötzsch (2014). For every entity name identified in the text, we query the Wikidata API to find possible matches. The mapped entities are then vectorized using a sentence-embedding model and stored in a vector database. To handle ambiguity, we keep the five most similar candidates for each extracted entity according to vector similarity.
To ensure consistency across triplets, the same LLaMa 3.3 70B Instruct model is used again to normalize entity and relation names. Given the list of candidate matches from the previous step, the model selects the most appropriate canonical form while taking the full sentence context into account. This process merges spelling variants and aliases referring to the same entity, resulting in a cleaner, unified graph structure.
Our goal is to build a graph that captures new information appearing in the latest news updates. Therefore, we discard any triplets that exactly match existing facts in Wikidata. Triplets absent from the knowledge base are treated as novel facts, as they are more likely to represent fresh events, and thus serve as valuable material for generating time-sensitive questions.
4.2 Question Types
The question generation stage begins with subgraph extraction from the constructed Knowledge Graph. We identify all subgraphs matching one of four predefined structural templates, each representing a distinct question type Yang et al. (2024):
Simple
relations: (Morty Smith | voice | Keisuke Chiba)
question: Who voiced Morty Smith?
answer: Keisuke Chiba
Set
question: What projects has Ryan Otter composed music for?
answer: Trigger, Method
Multi-Hop
answer: China
Conditional
question: Who performed at M-bar and met with Dmitry Dibrov?
answer: Roman Miroshnichenko
@@ 4.2 Question Types
Each selected subgraph is then passed to the language model, which generates a natural-language question–answer pair. The question is formulated as a fluent, contextually appropriate sentence, while the answer comprises one or more entities explicitly present in the subgraph.
4.3 QA Filtering
To ensure high-quality and contextually grounded question–answer pairs, we apply a multi-stage filtering pipeline combining linguistic validation, entity consistency, graph correspondence check, and LLM-based judgment.
1. Linguistic and Structural Filtering.
Firstly, we assess the grammatical correctness and fluency of each question using a RuRoBERTa-large model trained on the RuCoLa dataset555https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola Mikhailov et al. (2022). This step eliminates ungrammatical or poorly formed questions. Next, we perform Named Entity Recognition (NER) on the original source text using the Natasha library. The generated questions and answers are checked for the presence of these entities. Samples without explicit named entities are discarded to remove trivial, knowledge-free examples. We further filter out overly simplistic questions by evaluating them with smaller instruction-tuned LLMs (Qwen 2.5 7B Team (2024) and LLaMa 3 8B Grattafiori et al. (2024)) without context; if a model can answer a question directly from prior knowledge, it is excluded.
2. Graph Correspondence Filtering.
Each remaining QA pair is verified against the source subgraph used during question generation to ensure factual alignment. For every entity in the subgraph, we calculate the Levenshtein distance Levenshtein and others (1966) between its label and the text of both the question and answer. Each node in the graph (entity) is assigned 2 coefficients: question presence and answer presence. It is evaluated as the scaled Levenshtein distance between the name of the entity and the closest substring from the question and answer. These values allow us to check that all entities have been mentioned correctly.
3. LLM-as-Judge Evaluation.
In the final stage, we apply the LLM-as-Judge approach using POLLUX 7B Martynov et al. (2025), a model fine-tuned for fine-grained evaluation in Russian. Each QA pair is automatically rated along eight generative criteria: (1) Question literacy (grammar and style), (2) Clarity, (3) Naturalness, (4) Context sufficiency (answer can be found in the passage), (5) Context necessity (question depends on the passage), (6) Answer correctness, (7) Answer uniqueness, and (8) Answer literacy. More details on these criteria can be found in Appx. D.
Each criterion is transformed into a separate prompt with the specific scoring scale (0-2). For answer criteria (Literacy, Correctness, Uniqueness Based on Context), the prompt contains a news article, a question, and an answer; for other criteria, the answer is omitted. The example is classified as positive according to the particular criterion if the judge model assigns a rating of 1 or higher. This threshold was chosen to imitate the majority vote used in human evaluation.
| Criterion | Precision | Recall |
| Question Literacy | 0.96 | 0.99 |
| Question Clarity | 0.99 | 0.62 |
| Question Naturalness | 0.96 | 0.52 |
| Context Sufficiency | 0.94 | 0.71 |
| Context Necessity | 0.93 | 0.95 |
| Answer Correctness | 0.95 | 0.82 |
| Answer Uniqueness | 0.85 | 0.78 |
| based on context | ||
| Answer Literacy | 0.91 | 0.97 |
After all filtering stages, 150 high-quality questions per category are retained for the final benchmark dataset.
@@ 5 Experimental Setup
To construct our experimental RAG systems, we used the LangChain framework666https://pypi.org/project/langchain/. All texts from the Public Texts dataset are split into chunks of length 500 with an overlap of 100 characters. Each chunk is vectorized using the retrieval model of the evaluating RAG system with the corresponding document prefixes, and the resulting vectors are stored in a vector database.
During the search phase, we use the prompted retrieval model to find five of the most relevant texts that match the user’s query. Retrieved chunks are incorporated into a prompt provided to the LLM of the evaluated RAG system. If the total length of the filled-in prompt exceeds the model’s maximum context length, the contextual information is truncated to the required size. To accelerate LLM inference, we utilize the vLLM framework777https://github.com/vllm-project/vllm Kwon et al. (2023).
5.1 Experimental Setup Details
Embedding Model Prefixes
| Model | Query prefix | Text prefix |
| FRIDA | search_query: | search_document: |
LLM Prompt Template
‘‘Answer the question using the provided context.
Give me only an answer.
<context> {context} </context>
Question: {question}
Answer: ’’
Model Configuration
For serving models, we used the vLLM framework. The model parameters used are shown in Tab. 3.
| Model | TP | ML |
| Qwen 2.5 | 4 | 32768 |
| Qwen 2.5 | 1 | 32768 |
| Ruadapt Qwen | 4 | 32768 |
| Qwen 3 | 4 | 32768 |
| Gemma 3 | 1 | 131072 |
| Gemma 3 | 4 | 131072 |
Metrics
The performance of retrieval is measured by 3 metrics:
Hit Rate measures the proportion of queries for which the relevant document appears among the top-k retrieved results.
Mean Reciprocal Rank (MRR) evaluates ranking quality by measuring how highly the first relevant document is ranked, assigning higher scores when a relevant document appears earlier in the ranked list.
We evaluate End-to-end RAG systems with:
ROUGE-2 measures bigram overlap between the model output and the reference text, capturing local phrase-level similarity and rewarding matching adjacent word pairs.
ROUGE-L, measures the longest common subsequence between the model output and the reference text, capturing overlap in overall sentence structure.
The Judge Score is used to evaluate the overall answer quality, is calculated as the average of the automatic scores from Pollux 888https://ai-forever.github.io/POLLUX/ across multiple criteria (e.g., correctness, completeness, and relevance).